# On a generalisation of Cameron’s base size conjecture

### v2

@@ 1. Introduction
Acknowledgements. This paper is part of the author’s doctoral thesis under the supervision of Professor Tim Burness. The author thanks Professor Burness for his encouragement, feedback, and invaluable advice. She also thanks him for introducing her to this interesting problem. The author also thanks Frank Lübeck for sharing GAP-readable data on Foulkes functions of exceptional groups. Finally, she thanks the Heilbronn Institute for Mathematical Research for funding her PhD at the university of Bristol.
@@ 2. Preliminaries
In this section, we record some preliminary lemmas that we will repeatedly use to prove Theorem 1. We first discuss a small number theoretic lemma that we will repeatedly use to derive bounds on conjugacy class sizes, we then proceed to discuss probabilistic methods, and finally we briefly comment on computational methods.
Lemma 2.1.
The following are true:
  • If $a,b\in\mathbb{Z}^{+}$ with $a<b$, then
    $\frac{(q^{a}+1)}{(q^{b}+1)}<q^{a-b}$
This lemma is [51, 1.2 (i) and (iii)]
@@ 2.1. Probabilistic methods
Our main tool for proving our main result is the probabilistic method. Recall that if $G\leqslant{\rm Sym}(\Omega)$ is a transitive permutation group with point stabiliser $H$, then the fixed point ratio of $x\in G$ is defined as
${\rm fpr}(x,G/H)=\frac{|C_{\Omega}(x)|}{|\Omega|},$
where $C_{\Omega}(x)=\{\omega\in\Omega\,:\,\omega^{x}=\omega\}$ is the set of fixed points of $x$ and $x^{G}$ denotes the conjugacy class of $x$ in $G$. We will use a generalisation of the powerful probabilistic approach based on fixed point ratio estimates that was first introduced in [63] by Liebeck and Shalev for the study of bases, which is applicable in the more general regularity setting. We now state and prove some preliminary lemmas about fixed point ratios for finite almost simple permutation groups that will help us confirm the regularity of tuples.
We first state two lemmas that appeared in [2], which are natural generalisations of equivalent lemmas used for the study of base sizes of permutation groups, which can be found in [18]. Throughout the paper, if $n$ is a natural number, we will often write $[n]$ to denote the set $\{1,\ldots,n\}$.
@@
Let $G$ be a finite group and let $\tau=(H_{1},\ldots,H_{k})$ be a core-free tuple of subgroups of $G$. Suppose $x_{1},\ldots,x_{m}$ are elements in $G$ such that $|x_{i}^{G}|\geqslant B$ for all $i\in[m]$ and $\sum_{i=1}^{m}|x_{i}^{G}\cap H_{j}|\leqslant A_{j}$ for all $j\in[k]$. Then
@@
Let $G$ be an almost simple group with socle $G_{0}$, let $H$ be a core-free maximal subgroup of $G$, and let $x\in G$ be an element of prime order $r$. If $G_{1}=\langle G_{0},x\rangle$ and $H_{1}=H\cap G_{1}$, then ${\rm fpr}(x,G/H)={\rm fpr}(x,G_{1}/H_{1})$.
@@
First note that the maximality of $H$ implies that $G=G_{0}H$. In particular, $G=G_{1}H$ and therefore $G/H=G_{1}H/H$. This implies that $G_{1}$ also acts transitively on the cosets of $H$.
@@
Note that if $G_{0}\leqslant\langle G_{0},x\rangle\leqslant{\rm Inndiag}(G_{0})=G_{I}$, then Lemma 2.4 implies that
@@ 2.2. Computational methods
Let $G$ be a finite almost simple group of Lie type over $\mathbb{F}_{q}$ with socle $G_{0}$. There are two natural instances where we use Magma in our proofs. The first is in Section 4 where $G$ is classical and we aim to improve the existing fixed point ratio bound given in Theorem 3.9. There are some cases where we can only derive rather crude estimates, which prove the result when $q$ is sufficiently large, but do not suffice when $q$ is relatively small. In all of those cases, the dimension of the natural module for $G_{0}$ is relatively low (in particular no more than $8$), so we construct $G$ using the Magma function AutomorphismGroupSimpleGroup() and call MaximalSubgroups(G) to obtain a set of representatives of the $G$-classes of maximal subgroups of $G$. We then filter out the non-standard subgroups of $G$ and one can verify the necessary fixed point ratio bounds for all of those using the following functions:
@@
The function fpr takes as input a permutation group $G$, a subgroup $H$ of $G$ and an element $g\in G$ and computes ${\rm fpr}(g,G/H)$. The function MaxFpr takes as input a group $G$ a subgroup $H$ of $G$ and and returns the largest $t$ so that ${\rm fpr}(x,G/H)<|x^{G}|^{-t}$ for all $x\in G$ of prime order. The case where the computation using the code above is the most time consuming is when $G_{0}={\rm PSp}_{8}(3)$, in which case we can get output in less than an hour.
The second instance where we will be using Magma is in some cases where we want to check confirm the regularity of a non-standard tuple $\tau$ of $G$. As mentioned above, our standard method is to bound $\widehat{Q}(G,\tau)$ above by 11, but again, there are instances where the bounds we obtain are not good enough for small values of $q$, so in these cases we use the function RegTuplesPlus presented in [1, p.4] instead, to verify the regularity all $R_{\rm ns}(G)$-tuples of $G$. The function takes as input a finite group $G$, an ordered tuple $M=[H_{1},\ldots,H_{t}]$ of subgroups of $G$ and two positive integers $l$ and $N$ with $l\geqslant 2$. The output is the minimal integer $k\geqslant l$ such that all $k$-tuples of $G$ are regular, together with all the non-regular (k1)(k-1)-tuples of $G$. We refer the reader to [1] for the code and more information.
To be able to use RegTuplesPlus, both $|G|$ needs to be relatively small. In particular, we only make use of this function only if $G_{0}\in\{{}^{2}F_{4}(2)^{\prime},{}^{3}D_{4}(2)\}$ for exceptional groups and when $G$ is at most $8$-dimensional for classical groups. In the classical case, the maximal value of $q$ we use RegTuplesPlus for, highly depends on the dimension of the natural module and the Lie type of the group we are considering, but it is always relatively small. For example, if $G_{0}={\rm PSp}_{8}(q)$, then we only use Magma for the case $q=2$, whilst if $G_{0}=L_{2}(q)$, then we use Magma for $q\geqslant 31$.
@@ 3. Classical group preliminaries
Throughout this section, as well as Sections 4, 5 and 6, we will take $G$ to be a finite almost simple classical group over a finite field $\mathbb{F}_{q}$, where $q=p^{f}$ for some prime $p$ and some positive integer $f$. We will write $G_{0}$ for the socle of $G$ and $V$ will denote the natural module for $G_{0}$. Moreover, let $G_{I}$ denote the group ${\rm Inndiag}(G_{0})$ of inner-diagonal automorphisms of $G_{0}$.
@@
Let $G$ be a finite almost simple classical group. Then either $R_{\rm ns}(G)\leqslant 5$ with equality if and only if $G={\rm U}_{6}(2).2$. Moreover, one of the following holds:
  • (i)
    $\mathbb{P}(G,4)\rightarrow 1$ as $|G|\rightarrow\infty$;
  • (ii)
    $G_{0}={\rm Sp}_{6}(q)$ with $q$ even or $G_{0}=\Omega_{7}(q)$ with $q$ odd and $\mathbb{P}(G,6)\rightarrow 1$ as $|G|\rightarrow\infty$.
@@
    As we will explain in Section 3.3,we will be heavily using a function $\eta_{G}(t)$, which is a variant of a ’zeta type function’ introduced by Liebeck and Shalev in [61] in the proof of Theorem 3.1. We will heavily rely on results in [18] for this function, but as is the case with fixed point ratios, those results will not always be applicable in this more general setting, so we will need to refine those. We do this in Section 5.
@@ 3.1. Subgroup structure
The subgroup structure of finite classical groups is described by Aschbacher in [4], where he defines nine separate collections and proves that every maximal subgroup of a finite classical group is contained in one of those collections. Note that the situation is slightly different when ${\rm P\Omega}^{+}_{8}(q)$ or ${\rm PSp}_{4}(2^{f})$ and we discuss this below. The first eight collections, denoted by $\mathscr{C}_{1},\ldots,\mathscr{C}_{8}$ are often referred to as the geometric collections, and the final collection, which we denote by $\mathscr{S}$ is known as the non-geometric collection. The members of the geometric collections can be described in terms of the underlying geometry of $V$. For example, the members of $\mathscr{C}_{2}$ are stabilisers of appropriate direct sum decompositions, and similarly, the members of $\mathscr{C}_{4}$ and $\mathscr{C}_{7}$ are stabilisers of appropriate tensor product decompositions. The $\mathscr{S}$-collection consists of subgroups that correspond to some absolutely irreducible representation on $V$ of a covering group of $H\cap G_{0}$ (see [47, p.3] for a precise definition). Note that the groups in $\mathcal{S}$ are almost simple. Throughout this paper, we will adopt the definition of [47] for the collections of maximal subgroups, which differs slightly from the definition in the original paper by Aschbacher [4]. We give a brief description of each of the collections in Table 1.
Let $H$ be a maximal subgroup of $G$ such that $H\not\in\mathscr{S}$ and set $H_{0}=G_{0}\cap H$. If $H_{0}$ is maximal, then $H_{0}\in\mathscr{C}$, where $\mathscr{C}=\mathscr{C}_{1}\cup\cdots\cup\mathscr{C}_{8}$. If $G_{0}={\rm PSp}_{4}(2^{f})$, or ${\rm P\Omega}_{8}^{+}(q)$, then the existence of exceptional automorphisms gives rise to an extra collection of so-called novelty subgroups. Following the notation in [22], we will denote this class by $\mathscr{N}$. We give a description of all those subgroups in Table 2, although the only cases of groups in $\mathcal{N}$ that will be considered in this paper are the ones where $G_{0}={\rm PSp}_{4}(q)$ for even $q$ and $H$ is of type ${\rm O}^{\epsilon}_{2}(q)\wr S_{2}$ or ${\rm O}_{2}^{-}(q).2$. We note that Aschbacher’s original paper [4] only provides partial information for the case where $G_{0}={\rm P\Omega}^{+}_{8}(q)$ and $G$ contains triality graph automorphisms. In this case, the maximal subgroups were determined in [49] by Kleidman.
@@
In this paper we are interested in non-standard subgroup tuples of $G$. In the classical setting, a core-free subgroup of $G$ is non-standard if and only if it is not a subspace subgroup. We begin by giving a precise definition of a subspace subgroup.
@@
A core-free maximal subgroup $M$ of $G$ not containing $G_{0}$ is a subspace subgroup if one of the following holds:
@@
    $(G,M)$ is displayed in Table 3.
In Table 3 we list all the cases not covered by (i) and (ii) in Definition 3.4. Those are all instances where there exists some isomorphism $\phi:G\rightarrow K$ where ${\rm soc}(K)=T\cong G_{0}$ and one of the following holds:
@@
    $\phi(M)=L$ is a subspace subgroup of Type (i) or (ii) from Definition 3.4;
@@
    $T=A_{n}$ and $L$ does not act primitively on $\{1,\ldots,n\}$.
In the first column of Table 3 we list $G_{0}$, in the second column the type of $M$, in the third column we list the pairs with entries $T$ and the type of $L$, and finally, in the fourth column we mention the conditions under which such an isomorphism occurs, if any. We claim that Table 3 indeed covers all relevant cases. First suppose that $\phi$ is an automorphism. Noting that inner, diagonal, and field automorphisms do not permute the Aschbacher families, we deduce that $\phi$ must be a graph automorphism. Moreover, we must have $G_{0}\in\{{\rm P\Omega}_{8}^{+}(q),{\rm PSp}_{4}(q)^{\prime}\}$, as graph automorphisms of linear and unitary groups again preserve Aschbacher families. In those two cases, we verify that $\phi(M)$ is as claimed in Table 3.4 by inspecting [12, Tables 8.14, 8.50]. If $\phi$ is not an automorphism, then the possible isomorphisms of classical groups are listed in [47, 2.9.1], and by inspecting the Tables in [12, Chapter 8] and comparing orders of maximal subgroups, we verify that Table 3 is complete.
@@ 3.2. Conjugacy classes
In this section we will give a brief description of the conjugacy classes of prime order elements in classical groups, and we will establish the notation and terminology that we will be using throughout. We refer the reader to [22, Chapter 3] for a more detailed description of those classes and to [22, Chapter 2] for a more detailed analysis of the basic properties of the classical groups, as well as their diagonal, field, graph, and graph-field automorphisms. We note that due to a theorem of Steinberg [73, Theorem 30], every element in an almost simple classical group is a product of an inner, a diagonal, a field, and a graph automorphism.
@@
Here we will assume that $x$ is semisimple. We start by stating a fundamental theorem ([44, 4.2.2.(j)]) that we will repeatedly use throughout this paper:
@@
If $T$ is a finite simple group of Lie type and let $x\in T$ be a semisimple element. Then $x^{T}=x^{{\rm Inndiag}(T)}$.
@@
First assume that $c\geqslant 2$. Then [22, 3.1.3] implies that $x$ lifts to a unique element $\hat{x}$ of order $r$ in the corresponding matrix group. That is, $x=\hat{x}Z$, where $Z$ is the centre of this matrix group. We will now focus on the linear case, that is $G_{0}={\rm L}_{n}(q)$, to describe the structure of prime order classes, but one can get similar descriptions with minor modifications for the unitary, symplectic, and orthogonal groups.
Now note that $\hat{x}$ is diagonalisable in $\mathbb{F}_{q^{i}}$, but not any proper subfield, so Maschke’s Theorem implies that $\hat{x}$ fixes a direct sum decomposition
@@
is the $1$-eigenspace of $\hat{x}$ and $U_{j}$ is an $j$-dimensional subspace of $V$, on which $\hat{x}$ acts irreducibly.
@@
for some $\lambda_{j}\in\mathscr{S}_{r}$. We now note that the set of eigenvalues in $U_{j}\otimes\mathbb{F}_{q^{i}}$ coincides with some $\sigma$-orbit, so we may abuse notation and write
@@
With minor modifications we can get similar descriptions for odd order semisimple elements in all classical groups. For example if $G_{0}={\rm PSp}_{n}(q)$, then the main difference is that when $i$ is odd, then the $\sigma$-orbits come in inverse pairs $(\Lambda,\Lambda^{-1})$, and so every element of order $r$ is of the form $(\Lambda_{1},\Lambda_{1}^{-1})^{a_{1}},\ldots,(\Lambda_{t},\Lambda_{t}^{-1})^{a_{t}},I_{e}]$. And again, two elements of order $r$ are ${\rm Sp}_{n}(q)$-conjugate if and only if they have the same eigenvalues in $\mathbb{F}_{q^{i}}$, so we get a similar description for odd order semisimple classes. We also get a similar description of the centraliser structure of $\hat{x}$. In particular, if $\widehat{G}$ denotes the lift of $G_{I}$ in ${\rm GL}_{n}(q)$ (or ${\rm GL}_{n}(q^{2})$ if $G$ is unitary), then [22, Remark 3.3.4, Remark 3.4.4, Remark 3.5.6] give
@@
Finally, if $c=1$, then we still write $x=\hat{x}Z$, but in this case the lift in $\hat{x}$ is not unique, so additional conjugacy classes arise. A precise analysis of conjugacy in this case is given in [22, 3.2.2, 3.3.3] and the structure of $C_{G_{I}}(x)$ is recorded in [22, Tables B.3 and B.4].
@@
We finish this section by very briefly commenting on the semisimple involutions in ${\rm PGL}_{n}(q)$ and we refer the reader to [22, Sections 3.2.2, 3.3.2, 3.4.2, 3.5.2] for a detailed analysis of semisimple involutions in all classical almost simple groups of Lie type. So now let $G_{0}={\rm PGL}_{n}(q)$ and take $x$ to be a semisimple involution. For every $i$ such that $1\leqslant i\leqslant n/2$ we get a unique class of involutions with representative $t_{i}$, where $t_{i}=\hat{t}Z$ and $\hat{t}$ is an involution in ${\rm GL}_{n}(q)$ whose (1)(-1)-eigenspace has dimension $i$. If $x\in t_{i}^{G}$, then we will say that it is a $t_{i}$-involution. Finally, if $n$ is even, then we get an additional class of involutions with representative $t^{\prime}_{n/2}$. This class of involutions arises from the quadratic field extension $\mathbb{F}_{q^{2}}/\mathbb{F}_{q}$ and we direct the reader to [22, 3.2.2.2] for more details. The precise structure of $C_{G_{I}}(x)$ is recorded in [22, Table B.1].
@@
We now shift our attention to unipotent elements in finite classical groups. Let $x\in G$ be an element of order $p$. Then [22, 3.1.3] implies that $x$ again lifts to a unique element $\hat{x}$ of order $p$ in the corresponding matrix group, so that $x=\hat{x}Z$, where $Z$ again denotes the centre of this matrix group. We may now identify $\hat{x}$ with its Jordan decomposition on $V$ and write

### v3

@@ 1. Introduction
Acknowledgements. This paper is part of the author’s doctoral thesis under the supervision of Professor Tim Burness at the University of Bristol. The author thanks Professor Burness for his encouragement, feedback, and invaluable advice. She also thanks him for introducing her to this interesting problem. The author also thanks Frank Lübeck for sharing GAP-readable data on Foulkes functions of exceptional groups. Finally, she thanks the Heilbronn Institute for Mathematical Research for funding her PhD at the university of Bristol.
@@ 2. Preliminaries
In this section, we record some preliminary lemmas that we will repeatedly use to prove Theorem 2. We first discuss preliminaries related to probabilistic methods, and then we briefly comment on some of the main computational methods we apply in this paper.
@@ 2.1. Probabilistic methods
Our main tool for proving our main result is the probabilistic method. Recall that if $G\leqslant{\rm Sym}(\Omega)$ is a finite transitive permutation group with point stabiliser $H$, then the fixed point ratio of $x\in G$ is defined as
${\rm fpr}(x,G/H)=\frac{|C_{\Omega}(x)|}{|\Omega|}=\frac{|x^{G}\cap H|}{|x^{G}|}$
where $C_{\Omega}(x)=\{\omega\in\Omega\,:\,\omega^{x}=\omega\}$ is the set of fixed points of $x$ and $x^{G}$ denotes the conjugacy class of $x$ in $G$. We will use a generalisation of the powerful probabilistic approach based on fixed point ratio estimates that was first introduced in [64] by Liebeck and Shalev for the study of bases, which is also applicable in the more general regularity setting.
We first state two lemmas that appeared in [2, 2.1, 2.2], which are natural generalisations of equivalent lemmas used for the study of base sizes of permutation groups, which can be found in [18].
@@
Let $G$ be a finite group and let $\tau=(H_{1},\ldots,H_{k})$ be a core-free tuple of subgroups of $G$. Suppose $x_{1},\ldots,x_{m}\in G$ are such that $|x_{i}^{G}|\geqslant B$ for all $i\in[m]$ and $\sum_{i=1}^{m}|x_{i}^{G}\cap H_{j}|\leqslant A_{j}$ for all $j\in[k]$. Then
@@
Let $G$ be a primitive permuatation group with point stabiliser $H$, let $x\in G$, and let $1\neq N\trianglelefteqslant G$. If $G_{1}=\langle N,x\rangle$ and $H_{1}=H\cap G_{1}$, then ${\rm fpr}(x,G/H)={\rm fpr}(x,G_{1}/H_{1})$.
@@
First note that the maximality of $H$ implies that $G=NH$. In particular, $G=G_{1}H$ and therefore $G/H=G_{1}H/H$. This implies that $G_{1}$ also acts transitively on the cosets of $H$.
@@
Note that if $G$ is a finite almost simple group with socle $G_{0}$, $H$ is a maximal core-free subgroup of $G$ and $x\in G_{I}={\rm Inndiag}(G_{0})$, then $G_{0}\trianglelefteqslant\langle G_{0},x\rangle\leqslant G_{I}$, and Lemma 2.3 implies that
@@
We finally record a small number theoretic lemma that we will repeatedly use to derive fixed point ratio estimates:
Lemma 2.5.
The following are true:
  • If $a,b\in\mathbb{Z}^{+}$ with $b\leqslant a$, then
    $\frac{(q^{a}+1)}{(q^{b}+1)}<q^{a-b}$
Proof.
This lemma is [52, 1.2 (i) and (iii)]. ∎
@@ 2.2. Computational methods
Let $G$ be a finite almost simple group of Lie type over $\mathbb{F}_{q}$ with socle $G_{0}$. There are two natural instances where we use Magma in our proofs. The first is in Section 4 where $G$ is classical and we aim to improve the existing fixed point ratio bound given in Theorem 3.9. There are some cases where we can only derive rather crude estimates, which prove the result when $q$ is sufficiently large, but do not suffice when $q$ is relatively small. In the latter situation, the dimension of the natural module for $G_{0}$ is small (in particular no more than $8$), which means that we can construct $G$ using the Magma function AutomorphismGroupSimpleGroup() and call MaximalSubgroups(G) to obtain a set of representatives of the $G$-classes of maximal subgroups of $G$. We then filter out the non-standard subgroups of $G$ and one can verify the relevant fixed point ratio bounds using the following functions:
@@
The function fpr takes as input a permutation group $G$, a subgroup $H$ of $G$ and an element $g\in G$ and computes ${\rm fpr}(g,G/H)$. The function MaxFpr takes as input a group $G$ and a core-free subgroup $H$ of $G$ and returns the largest $t\in\mathbb{R}$ so that ${\rm fpr}(x,G/H)<|x^{G}|^{-t}$ for all $x\in G$ of prime order. The case where the computation using the code above is the most time consuming is when $G_{0}={\rm PSp}_{8}(3)$, in which case we can get output in less than an hour.
In some cases, we will also be using Magma to confirm the regularity of a non-standard tuple $\tau$ of $G$. As mentioned above, our standard method is to bound $\widehat{Q}(G,\tau)$ above by $1$, but again, there are instances where the bounds we obtain are not good enough for small values of $q$, so in these cases we use the function RegTuplesPlus (see [1, p.4]), to obtain $R_{\rm ns}(G)$. This takes as input a finite group $G$, an ordered tuple $M=[H_{1},\ldots,H_{t}]$ of subgroups of $G$ and two positive integers $l$ and $N$ with $l\geqslant 2$. The output is the minimal integer $k\geqslant l$ such that all $k$-tuples of $G$ containing subgroups in $M$ are regular, together with all the non-regular (k1)(k-1)-tuples of $G$. We refer the reader to [1] for the code and more information.
To be able to use RegTuplesPlus, $|G|$ needs to be relatively small. In particular, we only make use of this function if $G_{0}\in\{{}^{2}F_{4}(2)^{\prime},{}^{3}D_{4}(2)\}$ for exceptional groups and when $G$ is at most $8$-dimensional. In the classical case, the maximal value of $q$ we use RegTuplesPlus for, highly depends on the dimension of the natural module and the Lie type of the group we are considering, but it is always relatively small. For example, if $G_{0}={\rm PSp}_{8}(q)$, then we only use Magma for the cases $q=2,3$, whilst if $G_{0}={\rm L}_{2}(q)$, then we use Magma for $q\leqslant 31$.
@@ 3. Classical group preliminaries
Throughout this section, as well as Sections 4, 5 and 6, we will take $G$ to be a finite almost simple classical group over a finite field $\mathbb{F}_{q}$, where $q=p^{f}$ for some prime $p$ and some positive integer $f$. We will write $G_{0}$ for the socle of $G$ and $V$ will denote the natural module for $G_{0}$. Moreover, $G_{I}$ will denote the group ${\rm Inndiag}(G_{0})$ of inner-diagonal automorphisms of $G_{0}$.
@@
Let $G$ be a finite almost simple classical group. Then $R_{\rm ns}(G)\leqslant 5$ with equality if and only if $G={\rm U}_{6}(2).2$. Moreover $\mathbb{P}(G,4)\rightarrow 1$ as $|G|\rightarrow\infty$.
@@
    As we will explain in Section 3.3,we will be heavily using a function $\eta_{G}(t)$, which is a variant of a ‘zeta type function’ introduced by Liebeck and Shalev in [62] in the proof of Theorem 3.1. We will rely on results in [18] for this function, but as is the case with fixed point ratios, those results will not always be applicable in this more general setting, so we will need to refine those. We do this in Section 5.
@@ 3.1. Subgroup structure
The subgroup structure of finite classical groups is described by Aschbacher in [4], where he defines nine separate collections and proves that every maximal subgroup of a finite classical group is contained in one of those collections. Note that the situation is slightly different when ${\rm P\Omega}^{+}_{8}(q)$ or ${\rm PSp}_{4}(2^{f})$ and we discuss this below. The first eight collections, denoted by $\mathscr{C}_{1},\ldots,\mathscr{C}_{8}$ are often referred to as the geometric collections, and the final collection, which we denote by $\mathscr{S}$ is known as the non-geometric collection. The members of the geometric collections can be described in terms of the underlying geometry of $V$. For example, the members of $\mathscr{C}_{2}$ are stabilisers of appropriate direct sum decompositions, and similarly, the members of $\mathscr{C}_{4}$ and $\mathscr{C}_{7}$ are stabilisers of appropriate tensor product decompositions. The $\mathscr{S}$-collection consists of subgroups that correspond to some absolutely irreducible representation on $V$ of a covering group of $H\cap G_{0}$ (see [48, p.3] for a precise definition). Note that the groups in $\mathscr{S}$ are almost simple. Throughout this paper, we will adopt the definition of [48] for the collections of maximal subgroups, which differs slightly from the definition in the original paper by Aschbacher [4]. We give a brief description of each of the collections in Table 1.
Let $H$ be a maximal subgroup of $G$ such that $H\not\in\mathscr{S}$ and set $H_{0}=G_{0}\cap H$. If $H_{0}$ is maximal, then $H_{0}\in\mathscr{C}$, where $\mathscr{C}=\mathscr{C}_{1}\cup\cdots\cup\mathscr{C}_{8}$. If $G_{0}={\rm PSp}_{4}(2^{f})$, or ${\rm P\Omega}_{8}^{+}(q)$, then the existence of exceptional automorphisms gives rise to an extra collection of so-called novelty subgroups. Following the notation in [22], we will denote this class by $\mathscr{N}$. We give a description of all those subgroups in Table 2, although the only cases of groups in $\mathscr{N}$ that will be considered in this paper are the ones where $G_{0}={\rm PSp}_{4}(q)$ for even $q$ and $H$ is of type ${\rm O}^{\epsilon}_{2}(q)\wr S_{2}$ or ${\rm O}_{2}^{-}(q).2$. We note that Aschbacher’s original paper [4] only provides partial information for the case where $G_{0}={\rm P\Omega}^{+}_{8}(q)$ and $G$ contains triality graph automorphisms. In this case, the maximal subgroups were determined in [50] by Kleidman.
@@
In this paper we are interested in non-standard subgroup tuples of $G$. As we saw in the introduction, in the classical setting, a core-free subgroup of $G$ is non-standard if and only if it is not a subspace subgroup. We begin by giving a precise definition of a subspace subgroup.
@@
A core-free maximal subgroup $H$ of $G$ is a subspace subgroup if one of the following holds:
@@
    $(G,H)$ is displayed in Table 3.
In Table 3 we list all the cases not covered by (i) and (ii) in Definition 3.4. Those are all instances where there exists some isomorphism $\phi:G\rightarrow K$ where ${\rm soc}(K)=T\cong G_{0}$ and one of the following holds for $\phi(H)=L$:
@@
    $L$ is a subspace subgroup of $K$ of type (i) or (ii) from Definition 3.4;
@@
    $T=A_{m}$ and $L$ does not act primitively on $\{1,\ldots,m\}$.
In the first column of Table 3 we list $G_{0}$, in the second column the type of $H$, in the third column we list the pairs with entries $T$ and the type of $L$, and finally, in the fourth column we mention the conditions under which such an isomorphism occurs, if any.
We claim that Table 3 indeed covers all relevant cases. First suppose that $\phi$ is an automorphism. Noting that inner, diagonal, and field automorphisms do not permute the Aschbacher families, we deduce that $\phi$ must be a graph automorphism. Moreover, we must have $G_{0}\in\{{\rm P\Omega}_{8}^{+}(q),{\rm PSp}_{4}(q)^{\prime}\}$, as graph automorphisms of linear and unitary groups again preserve Aschbacher families. In those two cases, we verify that $\phi(H)$ is as claimed in Table 3.4 by inspecting the first two columns [12, Tables 8.14, 8.50].
If $\phi$ is not an automorphism, then the possible isomorphisms of classical groups are listed in [48, 2.9.1], and by inspecting the Tables in [12, Chapter 8] and comparing orders of maximal subgroups, we verify that Table 3 is complete.
@@ 3.2. Conjugacy classes
In this section we will give a brief description of the conjugacy classes of prime order elements in classical groups, and we will establish the notation and terminology that we will be using throughout. We refer the reader to [22, Chapter 3] for a more detailed description of those classes and to [22, Chapter 2] for a more detailed analysis of the basic properties of the classical groups, as well as their diagonal, field, graph, and graph-field automorphisms. We note that due to a theorem of Steinberg [76, Theorem 30], every element in an almost simple classical group is a product of an inner, a diagonal, a field, and a graph automorphism.
@@
Here we will assume that $x$ is semisimple. We start by stating a fundamental result ([45, 4.2.2.(j)]) that we will repeatedly use throughout this paper:
@@
If $T$ is a finite simple group of Lie type and let $x\in{\rm Inndiag}(T)$ be a semisimple element. Then $x^{T}=x^{{\rm Inndiag}(T)}$.
@@
First assume that $c\geqslant 2$ and let $\widehat{G}$ be defined according to Table 4 for each type of classical group. Then [22, 3.1.3] implies that $x$ lifts to a unique element $\hat{x}$ of order $r$ in $\widehat{G}$. That is, $x=\hat{x}Z$, where $Z$ is the centre of $\widehat{G}$.
We will now focus on the linear case, that is $G_{0}={\rm L}_{n}(q)$, to describe the structure of prime order classes. We will provide information on representatives, as well as the structure of centralisers. One can get similar descriptions with minor modifications for the unitary, symplectic, and orthogonal groups.
[…]
Table 4. Matrix groups of classical groups
Note that $\hat{x}$ is diagonalisable in $\mathbb{F}_{q^{i}}$, but not any proper subfield, so Maschke’s Theorem implies that $\hat{x}$ fixes a direct sum decomposition
@@
is the $1$-eigenspace of $\hat{x}$ and $U_{j}$ is an $i$-dimensional subspace of $V$, on which $\hat{x}$ acts irreducibly.
@@
for some $\lambda_{j}\in\mathscr{S}_{r}$. We now note that the set of eigenvalues of $\hat{x}$ on $U_{j}\otimes\mathbb{F}_{q^{i}}$ coincides with some $\sigma$-orbit, so we may abuse notation and write
@@
With minor modifications we can get similar descriptions for odd order semisimple elements in all classical groups. For example if $G_{0}={\rm PSp}_{n}(q)$, then the main difference is that when $i$ is odd, then the $\sigma$-orbits come in inverse pairs $(\Lambda,\Lambda^{-1})$, and so every element of order $r$ is of the form $(\Lambda_{1},\Lambda_{1}^{-1})^{a_{1}},\ldots,(\Lambda_{t},\Lambda_{t}^{-1})^{a_{t}},I_{e}]$. And again, two elements of order $r$ are ${\rm Sp}_{n}(q)$-conjugate if and only if they have the same eigenvalues in $\mathbb{F}_{q^{i}}$, so we get a similar description for odd order semisimple classes. We also get a similar description of the centraliser structure of $\hat{x}$. In particular, if $\widehat{G}$ is as defined in Table 4 for each type of classical group, then [22, Remark 3.3.4, Remark 3.4.4, Remark 3.5.6] give
@@
Finally, if $c=1$, then we still write $x=\hat{x}Z$, but in this case the lift in $\hat{x}$ might not be unique, so additional conjugacy classes arise when $r$ divides $n$. A precise analysis of conjugacy in this case is given in [22, 3.2.2, 3.3.3] and the structure of $C_{G_{I}}(x)$ is recorded in [22, Tables B.3 and B.4].
@@
We finish this section by very briefly commenting on the semisimple involutions in ${\rm PGL}_{n}(q)$ and we refer the reader to [22] for a detailed analysis of semisimple involutions in all classical almost simple groups. So now let $G_{0}={\rm PGL}_{n}(q)$ and take $x$ to be a semisimple involution. For every $i$ such that $1\leqslant i\leqslant n/2$ we get a unique class of involutions with representative $t_{i}$, where $t_{i}=\hat{t}Z$ and $\hat{t}$ is an involution in ${\rm GL}_{n}(q)$ whose (1)(-1)-eigenspace has dimension $i$. If $x\in t_{i}^{G}$, then we will say that it is a $t_{i}$-involution. Finally, if $n$ is even, then we get an additional class of involutions with representative $t^{\prime}_{n/2}$. This class of involutions arises from the quadratic field extension $\mathbb{F}_{q^{2}}/\mathbb{F}_{q}$ and we direct the reader to [22, 3.2.2.2] for more details. The precise structure of $C_{G_{I}}(x)$, as well as conditions for which $x\in G_{0}$ are recorded in [22, Table B.1].
@@
We now shift our attention to unipotent elements in finite classical groups. We refer the reader to [60] and [22] for more information. Let $x\in G$ be an element of order $p$. Then [22, 3.1.3] implies that $x$ again lifts to a unique element $\hat{x}$ of order $p$ in the corresponding matrix group, so that $x=\hat{x}Z$, where $Z$ again denotes the centre of this matrix group. We may now identify $\hat{x}$ with its Jordan decomposition on $V$ and write